In 15 years, Maya will be twice as old as David is now. In 15 years, David will be as old as Maya will be 10 years from now. How old is Maya now?
a.
20
c.
30
b.
25
d.
35

Answer by stanbon(75887)

In 15 years, Maya will be twice as old as David is now. In 15 years, David will be as old as Maya will be 10 years from now. How old is Maya now?
----
Equation(s)::
m+15 = 2d
d+15 = m+10
-----
Rearrange::
m - 2d = -15
m - d = 5
----------------
Subtract and solve for "d":
d = 20 (David's age now)
----
m = 5+d
m = 25 (Maya's age now)
=======================
Cheers,
Stan H.
-----------------------

a.
20
c.
30
b.
25
d.
35